Find the exact value of the expression. Do not use a calculator. 1 - cos2 69° - cos221° 


1 - cos2 69° - cos221° = _______ 

 (Simplify your answer, including any radicals. Use integers or fractions for any numbers in
Answer #1

SOLUTION :


1 - cos^2 (69) - cos^2 (21)

= sin^2 (69) - cos^2 (90-69)

= sin^2 (69) - sin^2 (69)

= 0 (ANSWER).


answered by: Tulsiram Garg
Answer #2

روکه - وی 2م) - / - برعهده) ) 2 (2) دم) - 2 sin(96-0) Coso = Since. ( ( - ) ك ) = 2 (ع) که)) - )= 2 (2 ) - 2 ( ره) - 2 و 2:S